# PENEX: AdaBoost-Inspired Neural Network Regularization

### v2

@@ 2.3 Relationship to Multi-Class AdaBoost
$f_{M}(\mathbf{x})\;=\;\sum_{m=1}^{M}\eta_{m}g_{m}(\mathbf{x}).$
These weak learners $g_{m}$, all together with their linear factors $\eta_{m}$, are fitted in a greedy fashion [^4] Also known as stagewise additive modeling using a multi-class exponential loss function (SAMME). according to the CONEX objective Sec. 2
@@
which ensures that the linear constraint $\sum_{j}f^{(j)}_{m}(\mathbf{x})=\sum_{j}\hat{\mathbf{y}}^{(j)}=0$ is fulfilled, for all $\mathbf{x}\in\mathcal{X}$ Sec. 2. In practice, the $g_{m}$ are typically represented by decision stumps, i.e., single-split decision trees (e.g., Hastie et al. (2017, p.362)). For completeness, we show the full AdaBoost algorithm in App. B.
@@
We now proceed to demonstrate formally that gradient descent on PENEX is analogous to AdaBoost in that gradient descent steps can be thought of as constrained parameters of weak learners that minimize the exponential loss. To see this, we consider a linearization about the previous parameters $\theta_{m-1}$, which assumes additivity as an approximation:
@@ 3 Related Work
The exponential loss is strongly associated with AdaBoost. Thus, besides discussing prior works on gradient-based exponential loss minimization and margin-maximization, we also cover related works in the field of boosting within deep learning.
@@
The idea of explicitly maximizing the classification margin in the context of deep learning has been explored in prior work: One line of work studies the effects of the spectral norm of the network’s Jacobian matrix (Sokolić et al., 2017) or network depth (Sun et al., 2016) on the classification margin to motivate regularization techniques. Elsayed et al. (2018) introduce a large‑margin objective that enforces margins at selected intermediate layers. The method adds several hyperparameters and requires computing per‑example gradient norms with respect to those activations (scaling with the number of layers and top‑k classes), making the method computationally demanding on deeper architectures. Another line of work by Jiang et al. (2019) shows that the distribution of the margin highly correlates with the generalization gap, which inspired Lyu et al. (2022) to develop a loss that penalizes deviations of the margin from its expected value.
@@
Tolstikhin et al. (2017); Grover & Ermon (2018); Giaquinto & Banerjee (2020) construct a “strong” generative model by greedily fitting multiple “weak” generative models in a boosting-like fashion. Similarly, but with classifiers instead of generative models, dense (Schwenk & Bengio, 1997) and convolutional neural networks (Drucker et al., 1992; LeCun et al., 1995; Taherkhani et al., 2020; Sun et al., 2021) have been used as weak learners for Boosting. In the context of language models, boosting has further been applied to prompt generation (Pitis et al., 2023; Zhang et al., 2024), text classification (Hou et al., 2023; Huang et al., 2020), text-summary based prediction (Manikandan et al., 2023) and machine teaching (Agrawal et al., 2024). All of these studies differ from the present work in what is considered as a “weak learner”. In the present work, this notion only arises implicitly (Sec. 2.3) and is given as the Jacobian, multiplied by a small parameter increment 10, instead of existing approaches that consider explicitly combining entire models to gradually build more complex ones.
@@ 4 Experiments
Figure 4: Performance Analysis on CIFAR-100. Larger means better. Results are computed from validation data. All hyperparameters have been tuned individually. PENEX (thick red) is an effective regularizer with often better generalization than other common regularization techniques (thin), and shows no signs of “overfitting” like cross-entropy training (orange).
@@ 4.1 Setup
In the main text, we provide a high-level description of the experimental setup. More details can be found in App. D.
@@
For fairness of comparison, we restrict ourselves to baselines that can be expressed as a single, architecture‑invariant loss function that does not require additional forward or backward passes. 1) Cross-entropy (CE). 2) Confidence penalty (Pereyra et al. (2017)), 3) Label smoothing (Szegedy et al. (2016)), and 4) Focal loss (Lin et al., 2018). The latter three are well-established regularization techniques.
@@
1-2) CIFAR-10/100 (Krizhevsky, 2009): Image classification with either $10$ or $100$ classes, using a convolutional neural network. 3) Noisy CIFAR-10: CIFAR-10 with 10% of training labels flipped uniformly at random. The test set does not have flipped labels. 4) Pathology Image Classification (PathMNIST; Kather et al. (2019); Yang et al. (2023)): Recognizing $9$ different tissue types from $10,000$ non-overlapping image patches from hematoxylin & eosin stained histological images, using a convolutional neural network. 5) ImageNet (Deng et al. (2009)): Recognizing $1000$ different classes from $\approx 14$ million images, using a vision transformer (Dosovitskiy et al., 2021). 6) News Classification (BBC News; Greene & Cunningham (2006)): Fine-tuning a pretrained RoBERTa-base language transformer encoder (Liu et al., 2019) for news article classification with $5$ classes.
@@
1) Accuracy (ACC): Proportion of correctly classified points. 2) Cross-Entropy (CE): Corresponds to $\mathcal{L}_{\mathrm{CE}}$, which (for classification) is equivalent to the negative log-likelihood. 3) Brier Score (BRIER; Brier (1950)): Mean squared difference between predicted class probabilities and one-hot encoded labels. 4) Expected Calibration Error (ECE; Guo et al. (2017)): Measures the discrepancy between a model’s predicted confidence and its actual accuracy. We change the sign in some metrics, such that a larger value indicates better performance for all metrics.
@@
Each method’s hyperparameters (except for the learning rate, which we keep fixed for all methods) are tuned on the validation set using the tree-structured Parzen estimator (e.g., Watanabe (2023)). We minimize cross-entropy loss over a fixed amount of epochs. For PENEX, we always estimate $\rho^{*}(\alpha)$ according to Alg. 1 (see App. C for details). Thus, PENEX only requires tuning the sensitivity $\alpha$. Tuning is performed individually for each experiment and method, except for ImageNet, where we use default parameters for each method (see LABEL:appx:imagenet_params).
Table 1: Test Set Performance. Larger means better, best is bold. Mean test result $\pm 1$ standard deviation after $200$ training epochs (for ImageNet: $300$ epochs). Standard deviation is obtained by running $100$ bootstrap evaluations (Efron & Tibshirani, 1993). Hyperparameters are tuned on validation data (except for ImageNet).
@@
4.2 Results
PENEX often achieves regularization performance that outperforms other regularization techniques (though taking a bit longer to converge), as can be seen in Fig. 4, showing multiple validation curves for CIFAR-100 (analogous plots for all others datasets are shown in LABEL:appx:all_validation_curves) as a function of the number of training epochs. We also provide test results in Tab. 1 that confirm the results from Fig. 4 on multiple datasets and domains. Remarkably, PENEX also outperforms other regularizers in the setting of artificial label noise. The only experiment where PENEX performs worse than other methods is the ImageNet experiment. However, ImageNet is a large data set and other regularization methods do also not benefit generalization much in this setting.
@@ 5 Discussion & Limitations
Conventional wisdom treats AdaBoost’s exponential loss as a flaw, inspiring variants that swap the exponential for the logistic loss (Friedman et al., 2000) and generalizations to arbitrary loss functions (Mason et al., 1999; Friedman, 2001). Our evidence suggests the opposite. PENEX typically outperforms standard cross-entropy training (and common regularization techniques; see Sec. 4.2). Remarkably, the resulting accuracy gap is wider than the well-documented (and usually modest) difference between AdaBoost and LogitBoost, two algorithms that typically achieve similar performance (e.g., Friedman et al., 2000). Explaining why the formulation introduced in the present work behaves so favourably in deep neural networks remains an open direction for future work.
@@
Convergence Speed & Instability.
For large values of $\alpha$ or large values of the learning rate, PENEX typically takes longer to converge and loss curves becomes more unsteady, compared to other methods (Fig. 4, LABEL:appx:parameter_analysis). However, PENEX typically achieves better generalization in the long run, for larger $\alpha$. We believe that adaptive scheduling of $\alpha$ could play an important role in optimizing training efficiency and stability, in addition to using smaller learning rates.
@@ 6 Conclusion
We introduced the penalized exponential loss (PENEX), a margin-promoting objective inspired by multi-class AdaBoost and designed for modern neural network training. Beyond its theoretical footing, PENEX reliably regularizes training and, in low-data settings, it typically surpasses existing methods. These results suggest PENEX is a strong default choice for practitioners and a promising basis for future work.
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. Yes
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. Not Applicable
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. Yes, anonymized source code, including instructions for optimal reproducibility, is included in the supplementary material.
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. Yes
    2. (b)
      Complete proofs of all theoretical results. Yes
    3. (c)
      Clear explanations of any assumptions. Yes
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). Yes
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). Yes
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). Yes
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). Yes
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. Yes
    2. (b)
      The license information of the assets, if applicable. Yes
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. Yes
    4. (d)
      Information about consent from data providers/curators. Not Applicable
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. Not Applicable
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. Not Applicable
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. Not Applicable
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. Not Applicable
PENEX: AdaBoost-Inspired
Neural Network Regularization
Supplementary Materials
@@ Appendix A Proofs
We consider the following blueprint:
  1. 1.
    We show that PENEX is Fisher consistent, conditionally on any $\mathbf{x}$ and for any $\rho>0$.
  2. 2.
    We show that PENEX is also unconditionally Fisher consistent.
Step 1
@@
Due to re-normalization, the factor $\rho/\alpha$ is canceled out and hence, the result follows.
Step 2
First, we note that the specific choice of $\rho$ does not matter for the result shown in Step 1, due to re-normalization. Consequently, we can pick a single, arbitrary $\rho>0$ that achieves Fisher consistency for all $\mathbf{x}$. We now consider the unconditional population equivalent of PENEX 3,
$\mathbb{E}\mathcal{L}_{\mathrm{\scriptscriptstyle PENEX}}(f)\;=\;\mathbb{E}_{\mathbf{x}}\left[\mathcal{L}_{\mathrm{\scriptscriptstyle PENEX}}^{y|\mathbf{x}}(f(\mathbf{x}))\right].$ (13)
It is easy to see that its minimizer is given by the function $f_{*}$ that minimizes the conditional PENEX $\mathcal{L}_{\mathrm{\scriptscriptstyle PENEX}}^{y|\mathbf{x}}(f(\mathbf{x}))$ for each $\mathbf{x}\in\text{supp}(P)$ 11. We consider an arbitrary function $f$ and see that
$\mathbb{E}[\mathcal{L}_{\mathrm{\scriptscriptstyle PENEX}}(f)]-\mathbb{E}[\mathcal{L}_{\mathrm{\scriptscriptstyle PENEX}}(f_{*})]=\int_{\mathbf{x}\in\text{supp}(P)}\left[\mathcal{L}_{\mathrm{\scriptscriptstyle PENEX}}^{y|\mathbf{x}}(f(\mathbf{x}))-\mathcal{L}_{\mathrm{\scriptscriptstyle PENEX}}^{y|\mathbf{x}}(f_{*}(\mathbf{x}))\right]p(\mathbf{x})d\mathbf{x}\geq 0,$
which concludes the proof. ∎
A.2 Proof for Obs. 2.1
Proof.
For simplicity, we assume that $\mathbf{x}$ can only take one single value, such that estimating P(Y|X)P(Y\,|\,X) simplifies to estimating the parameters of a multinoulli/categorical distribution, $P(y)$. Furthermore, we note that we make can make statements about the estimated $f$ by solving for $\hat{P}$ instead, since $\hat{P}(y)\propto\exp\left\{f^{(y)}\right\}$ constrains $f$ up to an additive constant. We consider a more explicit notation for the objective 4,
[…]
where H(PP^)H($\|\,\hat{P}$) denotes cross-entropy between $P$ and $\hat{P}$ and $\lambda>0$ is regularization strength. With $H(P)$ denoting Shannon entropy of $P$, we note that we can rewrite cross-entropy as
[…]
and see that the unique optimum for the trivial case $\lambda=0$ is given as $\hat{P}_{*}=P$, i.e., Fisher consistency holds. However, this is generally not the case if $\lambda>0$. To see this, we construct the Lagrangian
$\mathscr{L}(\hat{P})\;=\;\mathcal{L}(\hat{P})\;+\;\lambda^{\prime}\Big(\sum_{y=1}^{K}\hat{P}(y)-1\Big),$
with Lagrange multiplier $\lambda^{\prime}$. By the stationarity condition, it must hold for all $y$ that
[…]
At $\hat{P}=P$, 15 simplifies to
$\lambda\frac{\partial\Omega}{\partial\hat{P}(y)}\;=\;1-\lambda^{\prime},$
which means that $P$ can only be optimal if
$\frac{\partial\Omega}{\partial\hat{P}(y)}\Bigg|_{\hat{P}(y)=P(y)}=\text{const.},\;\forall y\in[K].$ (16)
[…]
$\frac{\partial\Omega}{\partial\hat{P}(y)}\;=\;1+\log(\hat{P}(y))$
[…]
A.3 Proof for Prop. 2.3
@@
For small $\eta$, the first constraint reduces to a half-space passing through the origin. This implies that the constraint can be strictly satisfied for $\eta=0$ and therefore, due to continuity, the feasible set in 10 (or equivalently 18) is nonempty for small enough $\eta$. We further note that $v(\eta)$, the minimum value in 18, is continuous for small $\eta$, see for example (Still, 2018, Ch. 5) and consider any sequence $\eta_{k}>0$ and corresponding $\Delta\theta_{mk}$, $k=1,2,\dots$ where $\eta_{k}\rightarrow 0$ and $\Delta\theta_{mk}$ belongs to the corresponding set of minimizers of 10. Our goal is to show that any limit point of the sequence belongs to the set of minimizer of 9 with $\eta=0$. To that extend, we first note that $\Delta\theta_{mk}$ remains bounded, since $\left\lVert\Delta\theta_{mk}\right\rVert_{2}^{2}=1$. Consider any limit point $\Delta\theta_{\infty}$ of $\Delta\theta_{mk}$ and pass to a subsequence $\Delta\theta_{mk}$ such that $\Delta\theta_{mk}\rightarrow\Delta\theta_{\infty}$. We now show that $\Delta\theta_{\infty}$ is a minimizer of 10 (or equivalently 18) for $\eta=0$. The limit point $\Delta\theta_{\infty}$ satisfies $\left\lVert\Delta\theta_{\infty}\right\rVert_{2}^{2}=1$ and the first constraint
@@
and is therefore feasible for the minimization in 18 when $\eta=0$. We conclude from the continuity of $v$ that $v(\eta_{k})\rightarrow v(0)$ and therefore $\Delta\theta_{\infty}$ is a minimizer of (10) for $\eta=0$. Hence the entire sequence $\Delta\theta_{mk}$ converges to the set of minimizers of (10) for $\eta=0$. As a result, $\Delta\theta_{\infty}$ satisfies the Karush-Kuhn-Tucker conditions corresponding to (18) for $\eta=0$. These can be expressed as
@@
where $\bar{\rho}\geq 0$ is related to a multiplier of the first constraint in 18. This means that $\Delta\theta_{\infty}\propto-\nabla_{\theta}\mathcal{L}_{\mathrm{\scriptscriptstyle PENEX}}(\theta_{m-1};\alpha)$ with $\rho=\bar{\rho}$, which concludes the proof.
A.4 Proof for Thm. 2.1
@@
Appendix B Stagewise Additive Modeling using a Multi-class Exponential loss

### v3

@@ 2.3 Relationship to Multi-Class AdaBoost
$f^{(y)}_{M}(\mathbf{x})\;=\;\sum_{m=1}^{M}\eta_{m}g^{(y)}_{m}(\mathbf{x}),$
for class $y\in[K]$. These weak learners $g_{m}$, all together with their linear factors $\eta_{m}$, are fitted in a greedy fashion [^3] Also known as stagewise additive modeling using a multi-class exponential loss function (SAMME). according to the CONEX objective Sec. 2
@@
which ensures that the linear constraint $\sum_{j}f^{(j)}_{m}(\mathbf{x})=\sum_{j}\hat{\mathbf{y}}^{(j)}=0$ is fulfilled, for all $\mathbf{x}\in\mathcal{X}$ Sec. 2. In practice, the $g_{m}$ are typically represented by decision stumps (e.g., Hastie et al. (2017, p.362)). For completeness, we show the full AdaBoost algorithm in App. C.
@@
We now proceed to demonstrate formally that gradient descent on PENEX is analogous to AdaBoost in that gradient descent steps on a neural network parameterization $f_{\theta}$ with parameters $\theta$ can be thought of as constrained parameters of weak learners that minimize the exponential loss. To see this, we consider a linearization about the previous parameters $\theta_{m-1}$, which assumes additivity as an approximation:
@@ 3 Related Work
The exponential loss is strongly associated with AdaBoost. Thus, besides discussing prior works on gradient-based exponential loss minimization and margin maximization, we cover related works on the intersection of boosting and deep learning.
@@
The idea of explicitly maximizing the classification margin in the context of deep learning has been explored in prior work: One line of work studies the effects of the spectral norm of the network’s Jacobian matrix (Sokolić et al., 2017) or network depth (Sun et al., 2016) on the classification margin to motivate regularization techniques. Elsayed et al. (2018) introduce a large‑margin objective that enforces margins at selected intermediate layers. The method adds several hyperparameters and requires computing per‑example gradient norms with respect to those activations (scaling with the number of layers and top‑k classes), making the method computationally demanding on deeper architectures. Another line of work by Jiang et al. (2019) shows that the distribution of the margin correlates with the generalization gap, which inspired Lyu et al. (2022) to develop a loss that penalizes deviations of the margin from its expected value.
@@
Tolstikhin et al. (2017); Grover & Ermon (2018); Giaquinto & Banerjee (2020) construct a “strong” generative model by greedily fitting multiple “weak” generative models in a boosting-like fashion. Similarly, but with classifiers instead of generative models, dense (Schwenk & Bengio, 1997) and convolutional neural networks (Drucker et al., 1992; LeCun et al., 1995; Taherkhani et al., 2020; Sun et al., 2021) have been used as weak learners for Boosting. In the context of language models, boosting has further been applied to prompt generation (Pitis et al., 2023; Zhang et al., 2024), text classification (Hou et al., 2023; Huang et al., 2020), text-summary based prediction (Manikandan et al., 2023) and machine teaching (Agrawal et al., 2024). All of these studies differ from the present work in what is considered as a “weak learner”. In the present work, this notion only arises implicitly (Sec. 2.3) and is given as the Jacobian, multiplied by a small parameter increment 10, instead of existing approaches that consider explicitly combining entire models (such as decision trees and neural networks) to gradually build more complex ones.
@@ 4 Experiments
Figure 4: Performance Analysis on CIFAR-100. Larger means better. Results are computed from validation data. All hyperparameters have been tuned individually. PENEX (thick red) is an effective regularizer with often better generalization than other common regularization techniques (thin), and shows no signs of “overfitting” like cross-entropy training (orange)
@@
Reproducibility.
We provide a supplementary .zip archive containing all code required to run the experiments and reproduce the figures and tables, along with a README.md containing instructions.
@@ 4.1 Setup
In the main text, we provide a high-level description of the experimental setup. More details can be found in App. E.
@@
For fairness of comparison, we restrict ourselves to baselines that can be expressed as a single, architecture‑invariant loss function that does not require additional forward or backward passes (more details in LABEL:appx:runtime). 1) Cross-entropy (CE). 2) Confidence penalty (Pereyra et al. (2017)), 3) Label smoothing (Szegedy et al. (2016)), and 4) Focal loss (Lin et al., 2018). The latter three are well-established regularization techniques.
@@
1-2) CIFAR-10/100 (Krizhevsky, 2009): Image classification with either $10$ or $100$ classes, using a convolutional neural network. An additional experiment with a residual neural network (He et al., 2016) is shown in LABEL:appx:alternative_architecture_experiment. 3) Noisy CIFAR-10: CIFAR-10 with 10% of training labels flipped uniformly at random. The test set does not have flipped labels. 4) Pathology Image Classification (PathMNIST; Kather et al. (2019); Yang et al. (2023)): Recognizing $9$ different tissue types from $10,000$ non-overlapping image patches from hematoxylin & eosin stained histological images, using a convolutional neural network. 5) ImageNet (Deng et al. (2009)): Recognizing $1000$ different classes from $\approx 14$ million images, using a vision transformer (Dosovitskiy et al., 2021). 6) News Classification (BBC News; Greene & Cunningham (2006)): Fine-tuning a pretrained RoBERTa-base language transformer encoder (Liu et al., 2019) for news article classification with $5$ classes.
@@
1) Accuracy (ACC): Proportion of correctly classified points ($1-\hat{\mathcal{R}}(f)$; see 6). 2) Cross-Entropy (CE): Corresponds to $\mathcal{L}_{\mathrm{CE}}$, which (for classification) is equivalent to the negative log-likelihood. 3) Brier Score (BRIER; Brier (1950)): Mean squared difference between predicted class probabilities and one-hot encoded labels. 4) Expected Calibration Error (ECE; Guo et al. (2017)): Measures the discrepancy between a model’s predicted confidence and its actual accuracy. We change the sign in some metrics, such that a larger value indicates better performance for all metrics.
@@
Each method’s hyperparameters (except for the learning rate, which we keep fixed for all methods) are tuned on the validation set using the tree-structured Parzen estimator (e.g., Watanabe (2023)). We minimize cross-entropy loss over a fixed amount of epochs. For PENEX, we always estimate $\rho^{*}(\alpha)$ according to Alg. 1 (see App. D for details). Thus, PENEX only requires tuning the sensitivity $\alpha$. Tuning is performed individually for each experiment and method, except for ImageNet, where we use default parameters for each method (see LABEL:appx:imagenet_params).
4.2 Results
PENEX often achieves regularization performance that outperforms other regularization techniques (though taking a bit longer to converge), as can be seen in Fig. 4, showing multiple validation curves for CIFAR-100 (analogous plots for all others datasets are shown in LABEL:appx:all_validation_curves) as a function of the number of training epochs. We also provide test results in Tab. 1 that confirm the results from Fig. 4 on multiple datasets, domains and metrics. Remarkably, PENEX also outperforms other regularizers in the setting of artificial label noise. The only experiment where PENEX performs worse than other methods is the ImageNet experiment. However, ImageNet is a large data set and other regularization methods do also not benefit generalization much in this setting.
Table 1: Test Set Performance. Larger means better, best is bold. Mean test result $\pm 1$ standard deviation after $200$ training epochs (for ImageNet: $300$ epochs). Standard deviation is obtained by running $100$ bootstrap evaluations (Efron & Tibshirani, 1993). Hyperparameters are tuned on validation data (except for ImageNet, see LABEL:appx:imagenet_params).
@@ 5 Discussion & Limitations
Conventional wisdom treats AdaBoost’s exponential loss as a flaw, inspiring variants that swap the exponential for the logistic loss (Friedman et al., 2000) and generalizations to arbitrary loss functions (Mason et al., 1999; Friedman, 2001). Our evidence suggests the opposite. PENEX typically outperforms standard cross-entropy training (and common regularization techniques; see Sec. 4.2). Remarkably, the resulting accuracy gap is wider than the well-documented (and usually modest) difference between AdaBoost and LogitBoost, two algorithms that typically achieve similar performance (e.g., Friedman et al., 2000). Explaining why the formulation introduced in the present work behaves so favorably in deep neural networks remains an open direction for future work.
@@
Instability.
For large values of $\alpha$ or large values of the learning rate, PENEX becomes more unsteady, compared to other methods (Fig. 4, LABEL:appx:parameter_analysis). We believe that adaptive scheduling of $\alpha$ could play an important role in optimizing training efficiency and stability, in addition to using smaller learning rates.
@@ 6 Conclusion
We introduced the penalized exponential loss (PENEX), a margin-promoting objective inspired by multi-class AdaBoost and designed for neural network training. We provided practical and theoretical motivation for its form and showed empirically that PENEX acts as a training regularizer: across various settings, it improves generalization over standard cross-entropy and, in low-data regimes, it often surpasses established alternatives. Beyond offering a practical loss function, our results also highlight a broader takeaway: the exponential loss can be useful beyond the setting of the AdaBoost algorithm. In particular, PENEX demonstrates that the margin-emphasizing behavior that underpins AdaBoost can be translated into modern neural optimization. Overall, we hope PENEX provides a simple and practical alternative to cross-entropy, particularly in limited-data regimes, and that it encourages further work on exponential-loss objectives beyond classical boosting.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Proofs
Due to re-normalization, the factor $\rho/\alpha$ is canceled out. Hence, for every $\mathbf{x}$ that has non-zero support, the true conditional distribution $P(y\,|\,\mathbf{x})$ is recovered. ∎
A.2 Proof for Prop. 2.3
@@
For small $\eta$, the first constraint reduces to a half-space passing through the origin. This implies that the constraint can be strictly satisfied for $\eta=0$ and therefore, due to continuity, the feasible set in 10 (or equivalently 14) is nonempty for small enough $\eta$. We further note that $v(\eta)$, the minimum value in 14, is continuous for small $\eta$, see for example (Still, 2018, Ch. 5) and consider any sequence $\eta_{k}>0$ and corresponding $\Delta\theta_{mk}$, $k=1,2,\dots$ where $\eta_{k}\rightarrow 0$ and $\Delta\theta_{mk}$ belongs to the corresponding set of minimizers of 10. Our goal is to show that any limit point of the sequence belongs to the set of minimizer of 9 with $\eta=0$. To that extend, we first note that $\Delta\theta_{mk}$ remains bounded, since $\left\lVert\Delta\theta_{mk}\right\rVert_{2}^{2}=1$. Consider any limit point $\Delta\theta_{\infty}$ of $\Delta\theta_{mk}$ and pass to a subsequence $\Delta\theta_{mk}$ such that $\Delta\theta_{mk}\rightarrow\Delta\theta_{\infty}$. We now show that $\Delta\theta_{\infty}$ is a minimizer of 10 (or equivalently 14) for $\eta=0$. The limit point $\Delta\theta_{\infty}$ satisfies $\left\lVert\Delta\theta_{\infty}\right\rVert_{2}^{2}=1$ and the first constraint
@@
and is therefore feasible for the minimization in 14 when $\eta=0$. We conclude from the continuity of $v$ that $v(\eta_{k})\rightarrow v(0)$ and therefore $\Delta\theta_{\infty}$ is a minimizer of (10) for $\eta=0$. Hence the entire sequence $\Delta\theta_{mk}$ converges to the set of minimizers of (10) for $\eta=0$. As a result, $\Delta\theta_{\infty}$ satisfies the Karush-Kuhn-Tucker conditions corresponding to (14) for $\eta=0$. These can be expressed as
@@
where $\bar{\rho}\geq 0$ is related to a multiplier of the first constraint in 14. This means that $\Delta\theta_{\infty}\propto-\nabla_{\theta}\mathcal{L}_{\mathrm{\scriptscriptstyle PENEX}}(\theta_{m-1};\alpha)$ with $\rho=\bar{\rho}$, which concludes the proof.
A.3 Proof for Thm. 2.1
@@
Appendix B Differences Between Thm. 2.1 and the Result by Schapire et al. (1998)
The earlier work by Schapire et al. (1998) shows that AdaBoost effectively increases margins. We recapitulate the Theorem 5 of Schapire et al. (1998), using the notation from the present work:
Theorem B.1 (Theorem 5 of Schapire et al. (1998)).
Consider a binary classification problem ($K=2$) with $f^{(2)}(\mathbf{x})\equiv-f^{(1)}(\mathbf{x})$ and assume that $f^{(1)}(\mathbf{x})\,=\,\sum_{m=1}^{M}\eta_{m}g_{m}(\mathbf{x})$ with weighted training errors
$\epsilon_{m}\,\coloneqq\,\mathbb{P}_{(\mathbf{x}_{i},y_{i})\sim\hat{P}_{m}}(\mathbbm{1}\{f^{(y_{i})}(\mathbf{x}_{i})<0\}),$
where the probability mass of the weighted empirical distribution $\hat{P}_{m}$ is given as
$\hat{P}_{m}(\mathbf{x}_{i},y_{i})\,\coloneqq\,\frac{\hat{P}_{m-1}(\mathbf{x}_{i},y_{i})\exp\{-\eta_{m}f_{m}^{(y_{i})}(\mathbf{x}_{i})\}}{Z_{m}}$ (16)
and
$\eta_{m}\coloneqq\frac{1}{2}\log\left(\frac{1-\epsilon_{m}}{\epsilon_{m}}\right)$ (17)
with uniform initialization $\hat{P}_{1}=\mathcal{U}([K])$. Then, for any $\gamma>0$, it holds that
$\hat{\mathbb{P}}(f_{M}^{(y)}(\mathbf{x})\,\leq\,\gamma)\,\leq\,2^{M}\prod_{m=1}^{M}\sqrt{\epsilon_{m}^{1-\gamma}(1-\epsilon_{m})^{1+\gamma}},$
for $f_{M}$ fitted via binary AdaBoost (Freund & Schapire, 1995).
A generalization of Thm. B.1 to multi-class AdaBoost has been provided by Wang (2013). These earlier results differ from our result (Thm. 2.1) in three fundamental ways [^5] We only refer to Thm. B.1 from hereon, because the same arguments can be made for the results presented by Wang (2013):
  1. 1.
    Thm. B.1 makes a strong structural assumption about the classifier and provides a bound in terms of the weighted training errors of the individual “intermediate learners” $f_{m}$. Our result, in contrast, makes no assumptions about $f$.
  2. 2.
    Thm. B.1 assumes that $f$ is fitted via the AdaBoost algorithm. Thm. 2.1 makes no assumption about how $f$ is trained.
  3. 3.
    We consider a new formulation of the exponential loss that differs from the standard exponential loss in that the class scores/logits $f^{(y)}$ do not need to sum up to zero (concretely, the $f^{(2)}(\mathbf{x})\equiv-f^{(1)}(\mathbf{x})$ assumption in Thm. B.1). Instead, PENEX incorporates a penalty to prevent the logits from diverging (see 3). The margin result from Thm. B.1 cannot be directly translated to the PENEX formulation.
Appendix C Stagewise Additive Modeling using a Multi-class Exponential loss